radomes.openskiesproject.org
Anik4 Anik4
  • 08-11-2015
  • Mathematics
Answered

If the probability of a win is 0.5 and the probability of a draw is 0.3, what is the probability of a loss?

Verified Answer :

Georgie24 Georgie24
  • 08-11-2015

the probability of loss is 0.2 because they must add up to 1.

Answer Link
averageeli averageeli
  • 08-11-2015

0.5+0.3=0.8
1-0.8=0.2
Probability of a loss is 0.2

Answer Link

Related Questions

Find the reciprocal of the number Multiply to check your answer 2 34 A 2 43 B 114 C 411 D 423
an alpha particle that is emitted as nuclear radiation has the same structure as a
a model car is one tenth of the size of the real car the models measures 42cm longs what the length of length of the real car
Please answer ASAP An explanation is a MUST REQUIREMENT in order to receive points and the Brainliest answer Thank you
Last week a pharmacy sold a bottle of shampoo for 500 This week the shampoo sells for 650 By what percent did the price increase
Add or subtract 14 74
What is the major atom in methane that contributes to climate change
8 What happened between the Danes and the Finns Summarize the story told by the court poet
Sherry Jones 59 years old a qualified chartered accountant decided to emigrate to the United Kingdom The move occurred at the end of the 2024 year of assessment
What is meant by the term mass movement